# GraphXForm: Graph transformer for computer-aided molecular design with application to extraction

### v1

@@ 3.4 Benchmark methods
We further use variational autoencoders (VAEs) that have been widely used as a generative model in ML-guided molecular design, cf. overviews in Refs.  (Sanchez-Lengeling & Aspuru-Guzik, 2018; Bilodeau et al., 2022; Anstine & Isayev, 2023). VAEs use an encoder-decoder structure: The encoder maps molecules to a continuous vector representation, referred to as latent space, from which the decoder reconstructs these molecules. The molecular latent space can facilitate exploration of the molecular space. That is, different optimization strategies can be employed to discover points in the latent space that correspond to promising novel molecules. These strategies include random sampling, Bayesian optimization, and GAs (Sanchez-Lengeling & Aspuru-Guzik, 2018; Rittig et al., 2023d; Anstine & Isayev, 2023).
We herein use the Junction Tree VAE (JT-VAE)  (Jin et al., 2018) in combination with GAs. The JT-VAE operates on molecular graphs and their non-cyclic abstractions (junction trees), and has shown a high rate of decoding vectors from its latent space to chemically valid molecules. For the solvent design task, we only consider molecules that conform to the alphabet $\Sigma$. We therefore train JT-VAE on a subset of the QM9 data set, which consists of about 128,000 molecules containing at most nine heavy atoms. We use GAs to optimize within the JT-VAE latent space because they enable exploring a large number of latent vectors, hence molecules, due to the linear scaling with the number of sampling points.
3.5 Design results
@@
The goal for all methods is to find suitable solvents with respect to the two objectives from Equations 7 and 9 for the two example problems, using the general setup outlined in Section 3.2.
@@
Table 1: Performance of different molecular design methods for the two example tasks. Each method is run across three different random seeds, with a time budget of 8 hours. We report the objective function evaluation of the best molecule found over all runs (‘max best’), best averaged over all the three runs $\pm$ standard deviation (‘mean best’), the average of the top 20 molecules of the best run (‘max top 20’), and the mean of the top 20 over all three runs $\pm$ standard deviation (‘mean top 20’).
@@
To contextualize the results of all methods, we selected all 6,098 compounds from the COSMObase 2020 database that conform to the alphabet $\Sigma$. We calculated the objective functions for those compounds that also meet the miscibility gap requirements, cf. (6). The top three compounds based on their objective values are shown in the first column in Figures 3 and 4. For the IBA task, the best molecule in this list achieves an objective value of 5.57, while for the TMB/DMBA task, the highest objective value is 3.03.
@@
Table 1 compares the performance of the different molecular design methods for the two solvent design tasks. We report the objective value of the best molecule found by each method, as well as the average objective values of the top 20 molecules. We also present results averaged over multiple runs with different random seeds, providing insight into the robustness of each method. Additionally, Figures 3 and 4 display the structural formulas of the top three molecules identified by each method.
@@
Figure 3: IBA task: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
@@
Figure 4: TMB/DMBA task: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
@@
In the IBA task, GraphXForm, REINVENT-Transformer, and Graph GA all identified the same best molecule, which has an objective value of 8.87. However, GraphXForm consistently found this molecule in every run, highlighting its stability. For example, in contrast, while Graph GA achieved the same average value for the top 20 molecules, its mean best objective was only 7.13.
@@
For the more challenging TMB/DMBA task, GraphXForm significantly outperforms all others across every metric. Notably, Graph GA performs better overall than the recent REINVENT-Transformer, likely due to REINVENT-Transformer’s sensitivity to initialization during RL fine-tuning. This is also reflected in its relatively high standard deviation compared to the other methods.
@@
Interestingly, the JT-VAE and STONED methods identified molecules with significantly lower objective values for the TMB/DMBA task when compared to other methods, with maximum values of approximately 2.16 and 2.39, respectively. However, their results for the IBA task (6.85 and 7.53) were closer to those of the other methods. We attribute this discrepancy to the nature of the tasks: The TMB/DMBA task is inherently more challenging and requires larger, more complex molecules, while the best-performing molecules in the IBA task were smaller. As JT-VAE was trained only on molecules with up to nine heavy atoms, its ability to explore larger molecules seems limited. Similarly, STONED struggled to effectively explore larger molecular structures.
@@
In summary, GraphXForm demonstrates highly promising results in both solvent design tasks, outperforming state-of-the-art molecular design methods in terms of identifying the best candidates and ensuring robustness in the design process. In the following sections, we further explore the flexibility of GraphXForm by imposing structural constraints on the designed molecules and starting the design process from initial molecular structures.
3.6 Including structural constraints in molecular design
@@
In addition to optimizing an objective, such as a physical property, it is often desirable to enforce specific structural characteristics in the generated molecular candidates. These structural constraints can include limitations on ring size or specific bonding patterns between atom types, which can improve chemical feasibility, such as enhancing synthesizability and stability or reducing toxic moieties. For instance, in the TMB/DMBA task, although the top molecules generated by GraphXForm exhibit high objective values, they possess several characteristics that pose challenges for solvent design. Notably, the presence of single nitrogen-nitrogen bonds often leads to instability (hyd, 2024), resulting in reactive compounds that are unsuitable for liquid-liquid extraction. Additionally, the existence of a nine-member ring could pose significant challenges to the solvent’s synthesizability (Steinborn et al., 2023).
As noted in Section 2.1, string-based molecular design methods cannot directly enforce such structural constraints. Typically, this is managed by assigning a very low objective value to molecules that violate these constraints. In contrast, by operating directly on the molecular graph, GraphXForm allows us to restrict the search space for many constraints by simply masking actions that would lead to constraint-violating molecules. To examine how structural constraints affect the performance of GraphXForm in finding molecules with high objective values, we impose the following constraints: We limit the ring size to a maximum of six atoms and disallow any bonds between two oxygen atoms as well as single bonds between two nitrogen atoms. We note that these constraints of course do not cover all possibilities for maintaining synthetic accessibility and chemical stability; instead, they serve as examples to illustrate the capabilities of GraphXForm.
@@
For the IBA task, the best molecule identified in Table 1 with objective value of 8.87 contains multiple nitrogen-nitrogen single bonds. We therefore rerun the design task with the additional structural constraints. The top three molecules designed by GraphXForm in these runs are shown in the top row of Figure 5. With constraints, we achieve a ‘mean best’ value of $8.23\pm 0.00$ and a ‘mean top 20’ value $7.98\pm 0.02$, which are only slightly lower than the values obtained without structural constraints. We also repeated these constrained design runs with REINVENT-Transformer, given its conceptual similarity to GraphXForm. However, REINVENT-Transformer only reached a ‘mean best’ value of $7.03\pm 1.14$, with its overall best molecule having an objective value of 7.87.
@@
The top three molecules identified by GraphXForm with the structural constraints for the TMB/DMBA task are shown in the bottom row of Figure 5. Under these constraints, GraphXForm achieves a ‘mean best’ value of $8.65\pm 0.00$ and a ‘mean top 20’ value of $8.42\pm 0.03$. Notably, these results remain superior to those of the other methods that do not impose structural constraints listed in Table 1.
@@
This further demonstrates GraphXForm’s flexibility and its ability to consistently design promising molecules.
Figure 5: Top three molecules designed by GraphXForm where: ring size is limited to a maximum of six atoms, any bond between oxygen atoms is disallowed, and single bonds between nitrogen atoms are disallowed.
3.7 Starting molecule design from initial structures
@@
Figure 6: Best molecule designed by GraphXForm when adding a maximum of 3, 5, and 7 atoms to a specific initial structure. We note that increasing the order of a bond within the initial structure is allowed. For TMB/DMBA, we also enforce that the designed molecule stays an alcohol.
@@
One advantage of GraphXForm, which is not feasible when constructing SMILES strings via next-token prediction due to the linear construction, is the ability to start the molecule design process from a predefined structure. This feature is particularly useful when a known candidate, potentially in practical use, possesses desirable properties but could benefit from slight modifications to improve the objective function. To demonstrate this, we consider the IBA task and begin with the best molecule from COSMObase 2020, which has an objective value of 5.57 (see Figure 4). We conduct three individual experiment runs, allowing GraphXForm to add up to 3, 5, and 7 atoms to the initial structure (and alter bond orders as needed). We also maintain the structural constraints outlined in Section 3.6. The original molecule, along with the best molecule from each of the three runs, is shown in the first row of Figure 6. In all cases, the designed molecules exhibit improvements in the objective value.
@@
We repeat a similar experiment for the TMB/DMBA task. During the list screening, we observed that the top three structures were all long-chain alcohols. However, long-chain alcohols tend to have melting points above room temperature, making them unsuitable as solvents for the processes under consideration. The melting point can be lowered by adding branches and increasing bond orders. Thus, we initiate three runs starting from the best molecule identified in the list screening, allowing the addition of up to 3, 5, or 7 atoms, respectively. Additionally, we constrain the design process to ensure the resulting molecule remains an alcohol by preventing any actions from being taken on the hydroxy group (one oxygen bonded to one hydrogen). The results are shown in the second row of Figure 6. While no better molecule is found when adding only up to 3 atoms, a notable improvement occurs when allowing to add 5 or 7 atoms.
This approach provides a powerful tool for situations where it is preferable not to start the design process from scratch but to make targeted modifications to existing molecules.
@@
We note that allowing the agent to remove atoms or bonds would not provide any benefit when starting the design from scratch, i.e., from a single atom. In fact, it would introduce additional redundancies into the search space. However, in the discussed scenarios where the design begins with an existing molecule, enabling the agent to remove atoms or bonds could in fact offer an advantage, as it allows for greater flexibility and deviations from the initial structure. We leave the exploration of this possibility for future work.
@@ Conclusion
We presented GraphXForm, a method for molecular design that follows the successful paradigm of self-supervised pretraining followed by (RL-based) fine-tuning, but operates directly on molecular graphs. By doing so, we addressed challenges faced by string-based methods, such as chemical validity or accomodating structural constraints. We introduced a technique derived from self-improvement learning to stably fine-tune a deep graph-transformer model with eight layers. On two solvent design tasks, we showed that GraphXForm can outperform state-of-the-art molecule design techniques. Additionally, our approach can flexibly adhere to specified structural constraints, such as bond types and functional groups, and can adaptively start the design process from existing molecular structures.
Looking ahead, several promising avenues for future development exist. First, we could expand GraphXForm to include more atoms and ionization states, as well as pretrain the network on much larger databases. We note that these features can already be integrated within our current framework; however, we restricted the set of possible atoms to ensure comparibility between our and other methods. We could further evaluate GraphXForm on additional solvent design tasks, as well as more extensively researched drug design tasks.
We also aim to incorporate more constraints to ensure the suitability of solvents for liquid-liquid separation processes, recognizing that numerous factors – such as boiling and melting points – are critical to a solvent’s effectiveness. This, however, will depend on the availability of reliable property predictors.
@@ Acknowledgements
This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 466417970 and 466387255 – within the Priority Programme “SPP 2331: Machine Learning in Chemical Engineering”. This work was performed as part of the Helmholtz School for Data Science in Life, Earth and Energy (HDS-LEE). Simulations were performed with computing resources granted by RWTH Aachen University under project “thes1232”. The authors gratefully acknowledge the Competence Center for Digital Agriculture (KoDA) at the University of Applied Sciences Weihenstephan-Triesdorf for providing additional computational resources.

### v2

@@ 3.3 Case study 2: Solvent design
JT-VAE adopts a VAE, a widely used generative model in ML-guided molecular design (Sanchez-Lengeling & Aspuru-Guzik, 2018; Bilodeau et al., 2022; Anstine & Isayev, 2023). VAEs use an encoder-decoder structure, where the encoder maps molecules into a continuous latent space, and the decoder reconstructs them from this representation. Particularly when the objective function is derived from a trained network, the molecular latent space can facilitate exploration of the molecular space. That is, different optimization strategies can be employed to discover points in the latent space that correspond to promising novel molecules. These strategies include random sampling, Bayesian optimization, and GAs (Sanchez-Lengeling & Aspuru-Guzik, 2018; Rittig et al., 2023d; Anstine & Isayev, 2023). In our work, we use JT-VAE in combination with GAs. JT-VAE operates on molecular graphs and their non-cyclic abstractions (junction trees), and it has demonstrated a high rate of decoding latent vectors into chemically valid molecules. Since we consider only molecules that conform to the alphabet $\Sigma$, we train JT-VAE on a subset of the QM9 dataset (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) consisting of approximately 128,000 molecules with at most nine heavy atoms.
3.3.1 Unconstrained results
@@
The goal for all methods is to find suitable solvents with respect to the two objectives from Equations 9 and 11 for the two example problems, using the general setup outlined in Section 3.3. In addition to the molecules shown in this section, a list of the top 20 molecules from the best runs can be found in Supplementary Figures 1-14.
@@
Table 2: Performance of different molecular design methods for the two example solvent design tasks. Each method is run across three different random seeds, with a maximum time budget of 8 hours. We report the objective function evaluation of the best molecule found over all runs (‘max best’), best averaged over all the three runs $\pm$ standard deviation (‘mean best’), the average of the top 20 molecules of the best run (‘max top 20’), and the mean of the top 20 over all three runs $\pm$ standard deviation (‘mean top 20’).
@@
To contextualize the results of all methods, we selected all 6,098 compounds from the COSMObase 2020 database that conform to the alphabet $\Sigma$. We calculated the objective functions for those compounds that also meet the miscibility gap requirements, see Eq. (8). The top three compounds based on their objective values are shown in the first column in Figures 3 and 4. For the IBA task, the best molecule in this list achieves an objective value of 5.57, while for the TMB/DMBA task, the highest objective value is 3.03.
@@
Table 2 compares the performance of the different molecular design methods for the two solvent design tasks. We report the objective value of the best molecule found by each method, as well as the average objective values of the top 20 molecules. We also present results averaged over multiple runs with different random seeds, providing insight into the robustness of each method. Additionally, Figures 3 and 4 display the structural formulas of the top three molecules identified by each method. We note that, in these results, we do not yet impose any structural constraints and focus solely on optimization.
@@
Figure 3: IBA task, unconstrained: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
@@
Figure 4: TMB/DMBA task, unconstrained: Top three molecules (with their corresponding SMILES string and objective value) identified by each method across all runs.
@@
In the IBA task, GraphXForm, REINVENT-Transformer, and Graph GA all identified the same best molecule, which has an objective value of 8.87. However, GraphXForm consistently found this molecule in every run, highlighting its stability. For example, in contrast, while Graph GA has a slightly higher average value for the best 20 molecules found, its mean best objective over all runs is only 7.13.
@@
For the more challenging TMB/DMBA task, GraphXForm outperforms all other methods across every metric. Notably, Graph GA outperforms REINVENT-Transformer overall, likely due to the latter’s sensitivity to initialization during RL fine-tuning. This factor is also reflected in its relatively high standard deviation. We also observe that the designs produced by Graph GA and GraphXForm exhibit substantial structural similarity, although GraphXForm makes additional refinements that lead to improved scores.
@@
Interestingly, the JT-VAE and STONED methods identified molecules with significantly lower objective values for the TMB/DMBA task when compared to other methods, with maximum values of approximately 2.16 and 2.39, respectively. However, their results for the IBA task (6.85 and 7.53) were closer to those of the other methods. We attribute this discrepancy to the nature of the tasks: The TMB/DMBA task is inherently more challenging and requires larger molecules with more complex branching, while the best-performing molecules in the IBA task were smaller. As JT-VAE was trained only on molecules with up to nine heavy atoms, its ability to explore larger molecules seems limited. Similarly, STONED struggled to effectively explore larger molecular structures.
@@
In summary, GraphXForm demonstrates highly promising results in both solvent design tasks, outperforming its comparison partners in terms of identifying the best candidates and ensuring robustness in the design process. In the following sections, we further explore the flexibility of GraphXForm by imposing structural constraints on the designed molecules and starting the design process from initial molecular structures.
3.3.2 Including structural constraints in molecular design
@@
In addition to optimizing an objective, such as a physical property, it is often desirable to enforce specific structural characteristics in the generated molecular candidates. These structural constraints can include limitations on ring size or specific bonding patterns between atom types, which can improve chemical feasibility, such as enhancing synthesizability and stability or reducing toxic moieties. For example, in the IBA task, although the top molecules generated by GraphXForm achieve high objective values, they exhibit features that pose challenges for solvent design. Notably, the presence of single nitrogen–nitrogen bonds and single nitrogen–carbon–nitrogen bonds may lead to instability in water (hyd, 2024), resulting in reactive compounds unsuitable for liquid–liquid extraction. Similarly, our top result for the TMB/DMBA task includes a carbon atom that is single-bonded to a nitrogen, a hydroxyl group, and a single hydrogen atom, which also raises concerns about instability.
A straightforward approach to enforcing such constraints is to assign an objective value of $\infty$- ∞ to any designed molecules that violate them, effectively discarding those candidates. In contrast, because GraphXForm operates directly on the molecular graph – transitioning from one molecule to the next without waiting for the completion of a string – we can flexibly restrict the search space by simply masking actions that would lead to constraint violations.
To examine how structural constraints affect the performance of GraphXForm in finding molecules with high objective values, we impose the following constraints: rings are limited to five or six atoms, and the following bonding patterns are disallowed: single bonds between two nitrogen atoms; single bonds between two oxygen atoms; a carbon atom single-bonded to two nitrogen atoms, unless the carbon is also double-bonded to an oxygen atom, forming a urea functional group; and a carbon atom single-bonded to a nitrogen atom, an oxygen atom, a hydrogen atom, and one other non-hydrogen functional group. We acknowledge that these constraints do not cover all aspects of synthetic accessibility, chemical stability, safety, or environmental impact. Rather, they serve as illustrative examples to demonstrate the capabilities of GraphXForm in incrementally limiting the designs to promising candidate structures that can be further evaluated by experts.
@@
For the IBA task, the top three molecules designed by GraphXForm under these constraints are shown in the top row of Figure 5. With constraints in place, we achieve a ‘mean best’ value of $7.28\pm 0.00$ and a ‘mean top 20’ value $6.85\pm 0.02$, which are only slightly lower than the values obtained without structural constraints. We also conducted constrained design runs with REINVENT-Transformer as a string-based counterpart; however, it reached only a ‘mean best’ value of $6.36\pm 0.39$, with its overall best molecule scoring $6.92$.
@@
The bottom row of Figure 5 displays the top three molecules identified by GraphXForm under the structural constraints for the TMB/DMBA task. Under these constraints, GraphXForm achieves a ‘mean best’ value of $8.54\pm 0.00$ and a ‘mean top 20’ value of $8.31\pm 0.06$. Notably, these results remain superior to those of the other methods without structural constraints, as reported in Table 2.
@@
These findings further demonstrate GraphXForm’s flexibility and its ability to consistently design promising molecules.
Figure 5: Top three molecules designed by GraphXForm under structural constraints on specific bonding patterns and ring sizes.
3.3.3 Starting molecule design from initial structures
@@
Figure 6: Best molecule designed by GraphXForm when augmenting a specific initial structure by adding up to 3, 5, or 7 atoms. Bonds can be added without restriction. For the TMB/DMBA task, the designed molecule is also required to remain an alcohol.
@@
One advantage of our molecular graph approach, which is not achievable out of the box with vanilla SMILES-based next-token prediction due to its linear construction, is the direct ability to initiate the design process from a predefined structure. This feature is especially useful when a known candidate, which already possesses desirable properties, could benefit from targeted modifications to further improve the objective function. To demonstrate this, we consider the IBA task and start with the best molecule from COSMObase 2020, which has an objective value of 5.57 (see Figure 4). We then conduct three separate experimental runs, allowing GraphXForm to add up to 3, 5, and 7 atoms to the initial structure while adding bonds as needed, all under the structural constraints outlined in Section 3.3.2. The original molecule and the best molecule from each of the three runs are shown in the top row of Figure 6. In every case, the same modified molecule is produced (with three additional atoms), resulting in an improved objective value.
@@
We repeat a similar experiment for the TMB/DMBA task. During list screening, we observed that the top three structures were all long-chain alcohols. However, long-chain alcohols tend to have melting points above room temperature, making them unsuitable as solvents for the intended processes. The melting point can be lowered by, e.g., adding branches. To address this, we again initiated three runs starting from the best molecule identified in the screening. These runs allowed the addition of up to 3, 5, and 7 atoms, thereby simulating branching of the long-chain alcohol. Additionally, we constrained the design process so that the resulting molecule remains an alcohol by preventing any modifications to the hydroxy group (an oxygen atom bonded to one hydrogen atom). The results, shown in the second row of Figure 6, indicate that while no improvement in the objective function is observed when adding only 3 atoms, a notable enhancement occurs when 5 or 7 atoms are added.
@@
This approach provides a powerful tool for cases where it is preferable to build upon an existing molecule rather than starting from scratch. We note that enabling the agent to remove atoms or bonds would not be beneficial when designing from a single atom, as it would introduce unnecessary redundancies into the search space. However, in scenarios where the design process begins with an existing molecule, allowing removal could offer additional flexibility and enable greater deviations from the initial structure. Extending the action space to include atom and bond removal is straightforward within our framework; however, we leave the exploration of this possibility for future work.
@@ Conclusion
We presented GraphXForm, a method for molecular design that follows the successful paradigm of self-supervised pretraining followed by (RL-based) fine-tuning, but operates directly on molecular graphs. By doing so, we addressed challenges faced by string-based methods, such as chemical validity or accomodating structural constraints. We introduced a technique derived from self-improvement learning to fine-tune a deep graph-transformer model. Based on the established drug design GuacaMol benchmark (Brown et al., 2019) and two solvent design tasks, we showed that GraphXForm can outperform state-of-the-art molecule design techniques. Additionally, our approach can flexibly adhere to specified structural constraints, such as bond types and functional groups, and can adaptively start the design process from existing molecular structures.
Looking ahead, several promising avenues for future development exist. First, we plan to expand the atom alphabet of GraphXForm and pretrain the network on significantly larger databases. Although these enhancements can be integrated into our current framework, we intentionally limited the atom set in this study to ensure comparability with other methods. Second, we aim to extend the action space by allowing the agent to remove bonds and atoms. While not necessary for our present investigation – and admittedly introducing further action symmetries – this addition would offer the model more possibilities when modifying starting molecules.
Additionally, future work could involve evaluating GraphXForm on a broader range of molecular design tasks and incorporating more design considerations. For instance, recent trends in ML-based drug development emphasize increased sample efficiency (Gao et al., 2022; Guo & Schwaller, 2024a, c; Rittig et al., ) and the integration of synthesizability models (Guo & Schwaller, 2024b; Gao et al., 2024). Although sample efficiency was not the primary focus of our current study, our method inherently maintains a kind of experience replay buffer, suggesting that techniques targeting replay buffers (Guo & Schwaller, 2024a, c) could be readily applied to further enhance GraphXForm. Furthermore, tweaking TASAR parameters (e.g., using a shallower beam width $\beta$ or a higher step size $\sigma$) and allowing more training batches in each fine-tuning epoch can help to make the method more sample-efficient.
We also aim to incorporate more constraints to ensure the suitability of the designed molecules for specific applications. In the case of solvents for liquid-liquid separation processes, this includes recognizing that numerous factors – such as boiling and melting points – are critical to a solvent’s effectiveness. This, however, will depend on the availability of reliable property predictors.
@@ Acknowledgements
This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 466417970 and 466387255 – within the Priority Programme “SPP 2331: Machine Learning in Chemical Engineering”. This work was performed as part of the Helmholtz School for Data Science in Life, Earth and Energy (HDS-LEE). Simulations were performed with computing resources granted by RWTH Aachen University under project “thes1232”. The authors gratefully acknowledge the Competence Center for Digital Agriculture (KoDA) at the University of Applied Sciences Weihenstephan-Triesdorf for providing additional computational resources. Furthermore, the authors thank Herbert Riepl for suggesting structural constraints.
Appendix A Atom alphabet
We list the atom alphabet for the different design tasks as a Python dictionary. We also provide the valence of an atom, as well as its formal_charge and chiral_tag – for atom specification in RDKit – if needed.
A.1 Drug design tasks
        "C": {"atomic_number": 6, "valence": 4},
        "C-": {"atomic_number": 6, "valence": 3, "formal_charge": -1},
        "C+": {"atomic_number": 6, "valence": 5, "formal_charge": 1},
        "C@": {"atomic_number": 6, "valence": 4, "chiral_tag": CHI_TETRAHEDRAL_CW},
        "C@@": {"atomic_number": 6, "valence": 4, "chiral_tag": CHI_TETRAHEDRAL_CCW},
        "N": {"atomic_number": 7, "valence": 3},
        "N-": {"atomic_number": 7, "valence": 2, "formal_charge": -1},
        "N+": {"atomic_number": 7, "valence": 4, "formal_charge": 1},
        "O": {"atomic_number": 8, "valence": 2},
        "O-": {"atomic_number": 8, "valence": 1, "formal_charge": -1},
        "O+": {"atomic_number": 8, "valence": 3, "formal_charge": 1},
        "F": {"atomic_number": 9, "valence": 1},
        "P": {"atomic_number": 15, "valence": 7},
        "P-": {"atomic_number": 15, "valence": 6, "formal_charge": -1},
        "P+": {"atomic_number": 15, "valence": 8, "formal_charge": 1},
        "S": {"atomic_number": 16, "valence": 6},
        "S-": {"atomic_number": 16, "valence": 5, "formal_charge": -1},
        "S+": {"atomic_number": 16, "valence": 7, "formal_charge": 1},
        "S@": {"atomic_number": 16, "valence": 6, "chiral_tag": CHI_TETRAHEDRAL_CW},
        "S@@": {"atomic_number": 16, "valence": 6, "chiral_tag": CHI_TETRAHEDRAL_CCW},
        "Cl": {"atomic_number": 17, "valence": 1},
        "Br": {"atomic_number": 35, "valence": 1},
        "I": {"atomic_number": 53, "valence": 1}
A.2 Solvent design tasks
        "C": {"atomic_number": 6, "valence": 4},
        "N": {"atomic_number": 7, "valence": 3},
        "O": {"atomic_number": 8, "valence": 2}
Appendix B Drug design results
Here, we present the complete results for the 20 goal-directed tasks of the GuacaMol benchmark. To put our results into even further perspective, we included additional baseline methods not mentioned in the main paper.
Table 3: GuacaMol goal-directed benchmarks
Benchmark SMILES GA‡ [1] SMILES LSTM‡ [2] Graph GA‡ [3] Reinvent¶ [4] GEGL¶ [5] CReM [6] MolRL-MGPT [7] GraphXForm (ours)
Celecoxib rediscovery 0.732 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Troglitazone rediscovery 0.515 1.000 1.000 1.000 0.552 1.000 1.000 1.000
Thiothixene rediscovery 0.598 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Aripiprazole similarity 0.834 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Albuterol similarity 0.907 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Mestranol similarity 0.790 1.000 1.000 1.000 1.000 1.000 1.000 1.000
C11H24 * 0.829 0.993 0.971 0.999 1.000 0.966 1.000 1.000
C9H10N2O2PF2Cl † 0.889 0.879 0.982 0.877 1.000 0.940 0.939 1.000
Median molecules 1 0.334 0.438 0.406 0.434 0.455 0.371 0.449 0.459
Median molecules 2 0.380 0.422 0.432 0.395 0.437 0.434 0.422 0.403
Osimertinib MPO 0.886 0.907 0.953 0.889 1.000 0.995 0.977 0.972
Fexofenadine MPO 0.931 0.959 0.998 1.000 1.000 1.000 1.000 0.999
Ranolazine MPO 0.881 0.855 0.920 0.895 0.933 0.969 0.939 0.944
Perindopril MPO 0.661 0.808 0.792 0.764 0.833 0.815 0.810 0.835
Amlodipine MPO 0.722 0.894 0.894 0.888 0.905 0.902 0.906 0.923
Sitagliptin MPO 0.689 0.545 0.891 0.539 0.749 0.763 0.823 0.965
Zaleplon MPO 0.413 0.669 0.754 0.590 0.763 0.770 0.790 0.727
Valsartan SMARTS 0.552 0.978 0.990 0.095 1.000 0.994 0.997 1.000
Deco hop 0.970 0.996 1.000 0.994 1.000 1.000 1.000 1.000
Scaffold hop 0.885 0.998 1.000 0.990 1.000 1.000 1.000 1.000
Total score 14.398 17.341 17.983 16.349 17.627 17.919 18.052 18.227
* Top-159, no stereochemistry. † Top-250, no stereochemistry. ‡ Results as reported in [8]. ¶ Results as reported in [7].
[1] Naruki Yoshikawa, Kei Terayama, Masato Sumita, Teruki Homma, Kenta Oono, and Koji Tsuda. Population-based de novo molecule generation, using grammatical evolution. Chemistry Letters, 47(11):1431–1434, 2018.
[2] Marwin HS Segler, Thierry Kogej, Christian Tyrchan, and Mark P Waller. Generating focused molecule libraries for drug discovery with recurrent neural networks. ACS central science, 4(1):120–131, 2018.
[3] Jan H Jensen. A graph-based genetic algorithm and generative model/monte carlo tree search for the exploration of chemical space. Chemical science, 10(12):3567–3572, 2019.
[4] Marcus Olivecrona, Thomas Blaschke, Ola Engkvist, and Hongming Chen. Molecular de-novo design through deep reinforcement learning. Journal of Cheminformatics, 9, 2017.
[5] Sungsoo Ahn, Junsu Kim, Hankook Lee, and Jinwoo Shin. Guiding deep molecular optimization with genetic exploration. Advances in Neural Information Processing Systems, 33, 12008–12021, 2020.
[6] Pavel Polishchuk. CReM: chemically reasonable mutations framework for structure generation. Journal of Cheminformatics, 12, 2020.
[7] Xiuyuan Hu, Guoqing Liu, Yang Zhao, and Hao Zhang. De novo drug design using reinforcement learning with multiple gpt agents. Advances in Neural Information Processing Systems, 36, 7405-7418, 2023.
[8] Brown, N.; Fiscato, M.; Segler, M. H. S.; Vaucher, A. C. GuacaMol: Benchmarking Models for de Novo Molecular Design. Journal of Chemical Information and Modeling, 59(3):1096–1108, 2019.
Appendix C Solvent design results
C.1 GraphXForm
Figure 7: Top 20 molecules: GraphXForm, IBA task
Figure 8: Top 20 molecules: GraphXForm, TMB/DMBA task
C.2 List screen
Figure 9: Top 20 molecules: List screen, IBA task
Figure 10: Top 20 molecules: List screen, TMB/DMBA task
C.3 JT-VAE
Figure 11: Top 20 molecules: JT-VAE, IBA task
Figure 12: Top 20 molecules: JT-VAE, TMB/DMBA task
C.4 STONED
Figure 13: Top 20 molecules: STONED, IBA task
Figure 14: Top 20 molecules: STONED, TMB/DMBA task
C.5 Graph GA
Figure 15: Top 20 molecules: Graph GA, IBA task
Figure 16: Top 20 molecules: Graph GA, TMB/DMBA task
C.6 REINVENT-Transformer
Figure 17: Top 20 molecules: REINVENT-Transformer, IBA task
Figure 18: Top 20 molecules: REINVENT-Transformer, TMB/DMBA task
C.7 GraphXForm: Structural constraints
Figure 19: Top 20 molecules: GraphXForm, IBA task with structural constraints
Figure 20: Top 20 molecules: GraphXForm, TMB/DMBA task with structural constraints